A boy rides away from home in an automobile at a rate of 28 miles per hour and walks back at a rate of 4 miles an hour. The round trip requires 2 hours. How far does he ride?
Found 2 solutions by greenestamps, Alan3354:
Answer by greenestamps(13196)



Let x be the number of hours he rides away from home at 28 mph.

Then, since the speed of the automobile is 7 times his walking speed, the number of hours he spends walking is 7x.

Then the total time, which is 2 hours, is equal to x plus 7x, or 8x.

8x = 2 --> x = 1/4

So the amount of time he spends in the automobile is x = 1/4 hour.

The distance he travels in 1/4 hour at 28 mph is 7 miles.

Answer: 7 miles

Answer by Alan3354(69443)

A boy rides away from home in an automobile at a rate of 28 miles per hour and walks back at a rate of 4 miles an hour. The round trip requires 2 hours. How far does he ride?
---------
Avg speed of the round trip = 2*28*4/(28+4) = 7 mi/hr
RT = 7*2 = 14 miles
--> 7 miles each way
=============
The formula is similar to parallel flows, parallel work, parallel resistance, etc.
Just a factor of 2 different.
----
I'm not a proponent of memorizing formulas in general, but this one is simple and similar to others.